# RRWNet: Recursive Refinement Network for Effective Retinal Artery/Vein Segmentation and Classification

### v2

@@ 1 Introduction
The characteristics of retinal blood vessels (BV), including their caliber and configuration, serve as valuable biomarkers for diagnosing and monitoring several diseases and medical conditions, such as glaucoma, age-related macular degeneration (AMD), diabetic retinopathy (DR), and hypertension Sun et al., (2009); Abràmoff et al., (2010); Kanski and Bowling, (2011). These alterations can be readily identified by trained ophthalmologists through analysis of color fundus images acquired via retinography, a non-invasive and cost-effective imaging technique that involves capturing photographs of the retina through the dilated pupil. By virtue of its affordability and lack of invasiveness, retinography has become widely adopted in clinical practice, research investigations, and population-wide screening programs. An example of a retinography image is shown in Fig. 1 (left).
@@
A comprehensive analysis of the retinal vasculature requires the segmentation of blood vessels and their subsequent classification into arteries and veins (A/V). This process yields separate A/V segmentation maps (as illustrated in Fig. 1, right), enabling the quantification of various diagnostically relevant vessel characteristics, such as width, diameter, and tortuosity. Furthermore, accurate measurement of these characteristics facilitates the calculation of more complex biomarkers, including the arteriolar-to-venular diameter ratio (AVR) Ikram et al., (2004); Sun et al., (2009); Hatanaka et al., (2005).
However, manual execution of these tasks is inherently laborious, leading to increased costs, and is prone to human error, which negatively impacts both reproducibility and quality of care. To address these limitations, several automated approaches have been proposed for the simultaneous segmentation and classification of arteries and veins Mookiah et al., (2021).
Current state-of-the-art methods predominantly leverage fully convolutional neural networks (FCNNs) Long et al., (2015) for this purpose Galdran et al., (2019); Hemelings et al., (2019); Morano et al., (2021); Chen et al., (2022); Galdran et al., (2022); Karlsson and Hardarson, (2022); Hu et al., (2024). Most prevalent approaches classify each pixel into one of four classes: background, artery, vein, and crossing (representing regions where arteries and veins overlap). Additionally, some methods incorporate an “uncertain” class to account for pixels presenting ambiguous characteristics Galdran et al., (2019); Hemelings et al., (2019); Morano et al., (2021); Chen et al., (2022); Galdran et al., (2022); Karlsson and Hardarson, (2022); Hu et al., (2024). Conversely, some recent approaches Chen et al., (2022); Morano et al., (2021) frame the A/V segmentation and classification task as a multi-label segmentation problem. This framework entails training the network to simultaneously segment arteries, veins, and BV (i.e., both arteries and veins) as separate classes, allowing a single pixel to be assigned to one or more classes.
@@
Figure 2: Examples of manifest classification errors produced by a state-of-the-art FCNN-based method Morano et al., (2021). (1-3) While most of the vessel is classified as artery, the model misclassifies the last part as vein. (4-6) The model often confuses the classification of vessels in crossing areas. These errors are easily detected by a human observer because they are inconsistent with the overall structure of the vascular tree, hence the term “manifest”.
These errors arise from the propensity of FCNN-based models to classify vessels based on local characteristics of the input image, neglecting the global structural context of the vascular tree. To mitigate these errors, several approaches have been proposed Girard et al., (2019); Kang et al., (2020); Chen et al., (2022); Hu et al., (2024). Some methods Girard et al., (2019); Kang et al., (2020) employ ad hoc post-processing techniques based on graph propagation. Alternatively, other methods Chen et al., (2022); Hu et al., (2024) have proposed combining standard pixel-wise segmentation losses with adversarial losses Goodfellow et al., (2014) and custom-designed losses focused on specific characteristics of the predicted maps, such as topological consistency. This combined approach aims to guide the model towards generating more topologically consistent segmentations. Despite their contributions, these approaches exhibit limitations in their applicability beyond the specific task of blood vessel segmentation and classification (i.e., limited generalizability), and their overall effectiveness in mitigating manifest classification errors remains limited.
@@ 2.1 Vessel segmentation and classification
The first methods for vessel segmentation on color fundus images were based on ad hoc image processing techniques Jiang and Mojon, (2003); Nain et al., (2004); Tolias and Panas, (1998); Staal et al., (2004) or traditional learning models such as artificial neural networks Sinthanayothin et al., (1999); Marín et al., (2011). Today, FCNNs based on the U-Net architecture Ronneberger et al., (2015) have become the state of the art Jiang et al., (2018); Oliveira et al., (2018); Jin et al., (2019); Wang et al., (2021); Liu et al., 2023c ; Liu et al., 2023a ; Liu et al., 2023b .
Until recently, A/V segmentation and classification was treated as a two-step process, where A/V classification was performed only on pixels previously identified as BV by a segmentation algorithm Relan et al., (2014); Dashtbozorg et al., (2014); Zamperini et al., (2012); Estrada et al., (2015); Welikala et al., (2017). In addition, many works restricted the classification to small regions, usually around the optic disc Zamperini et al., (2012); Relan et al., (2014). The first work that deals with the whole vascular tree Dashtbozorg et al., (2014) proposed a graph-based method that takes as input a previously segmented vascular graph and the input image and obtains two separate graphs for arteries and veins. Later, Welikala et al., (2017) were the first to propose the use of a CNN for the classification stage. Although these methods achieved reasonable performance, they were limited by the quality of the initial vessel segmentation.
To avoid this problem, several works have addressed the simultaneous segmentation and classification of retinal vessels either as a semantic segmentation task of three to four classes (background, artery, vein, uncertain) Xu et al., (2018); Girard et al., (2019); Hemelings et al., (2019); Galdran et al., (2019); Ma et al., (2019); Kang et al., (2020); Morano et al., (2021); Chen et al., (2022); Karlsson and Hardarson, (2022); Galdran et al., (2022) or as a multi-label segmentation task with multiple targets Morano et al., (2021); Chen et al., (2022) (arteries, veins, blood vessels). The latter approach has the advantage of providing continuous and thus more topologically consistent segmentation maps of arteries and veins, since vessel crossings are considered as both classes at the same time.
@@ 2.2 Reducing manifest classification errors
Ad hoc post-processing methods Girard et al., (2019); Kang et al., (2020) typically involve graph-based operations on the vessel graph extracted from previously segmented maps. These methods usually require additional input information (such as the location of the optic disc), and their effectiveness is severely limited by the quality of the initial segmentation maps, which hinders their generalization capabilities.
Conversely, learning-based methods Hu et al., (2024); Chen et al., (2022); Karlsson and Hardarson, (2022) aim to address manifest classification errors by incorporating additional losses and architectural modifications. Hu et al., (2024) introduced a multi-class point consistency module to generate artery and vein skeletons, which are then used to compute different consistency losses aimed at improving topological consistency and mitigating classification errors. Similarly, Chen et al., (2022) proposed a GAN-based method with a topological loss. In particular, they proposed to use a discriminator designed to rank, from lowest to highest, the topological connectivity of the ground truth, the predicted mask, and a randomly transformed mask. The ranking error is used as a loss to encourage the model to produce more topologically consistent segmentation maps. In addition, they proposed a new module that extracts the high-level topological features of the images to force the model to predict vascular segmentation maps with a topology similar to that of the manual annotations. Although these methods are interesting and fairly effective, they rely on task-specific mechanisms, and their performance remains limited.
@@ 2.3 Iterative refinement
In recent years, several iterative refinement approaches have been proposed to improve segmentation performance in both natural and medical image analysis Pinheiro and Collobert, (2014); Sironi et al., (2016); Newell et al., (2016); Shen et al., (2017); Januszewski et al., (2016); Mosinska et al., (2018); Karlsson and Hardarson, (2022); Galdran et al., (2022). These methods use an iterative prediction process via a classifier that receives as input the result from the previous iteration(s) and, optionally, the input image, addressing the errors made in earlier iterations.
@@
A common approach consists of stacking multiple deep modules and training them in an end-to-end fashion Newell et al., (2016); Shen et al., (2017); Karlsson and Hardarson, (2022); Galdran et al., (2022). This allows for modules specialized in solving the errors of the previous modules. For example, Newell et al., (2016) proposed a novel architecture composed of eight consecutive modules for pose estimation in natural images. During training, all modules receive supervision through comparison of their outputs with the ground truth. However, such methods often require a substantial number of parameters, leading to high memory and computational costs during both training and inference. To address this limitation, recent work has proposed the use of very lightweight encoder-decoder networks Karlsson and Hardarson, (2022); Galdran et al., (2022). Using a custom architecture consisting of four U-Net-like networks, Karlsson and Hardarson, (2022) achieved state-of-the-art performance on various A/V segmentation and classification benchmarks. However, their approach requires extensive hyperparameter tuning to perform well, and the optimal hyperparameters were found to be dataset-specific. Their final model was achieved through an exhaustive search exploring various network configurations and loss functions. In particular, the authors experimented with different numbers of networks, layers, levels, and kernels in each network/layer, as well as different weights for the loss terms and the regularization parameters. Additionally, the generalization capabilities of the method have not been thoroughly evaluated. Galdran et al., (2022) proposed a similar approach for the same task using only two stacked U-Net-like networks, although with limited performance.
@@
An alternative approach consists of refining the predictions using a single recursive network Pinheiro and Collobert, (2014); Mosinska et al., (2018). While the memory requirements during training of this approach remain constant, as the gradients for each iteration must be stored, the number of parameters is much lower, which makes it more efficient at test time. In this line of work, Pinheiro and Collobert, (2014) proposed to perform semantic segmentation of natural images by using a CNN network that subsequently refines the predicted maps at different scales. This approach is focused on increasing the spatial context of the network, so that it models non-local dependencies (of higher level) in the scenes. In this way, the authors manage to make the network give rise to more structurally coherent predictions. The problem with this method is that the CNN is applied pixel-wise, making it very inefficient in terms of computational cost. This problem is addressed by Mosinska et al., (2018) by using a FCNN at a constant scale. Their method involves training a FCNN that recursively refines an initial segmentation over multiple iterations. In the first iteration, the network receives the input image and an empty segmentation map (all zeros). In subsequent iterations $k\in\{1,...,K\}$, it receives the same image together with the segmentation map obtained in iteration $k-1$. The training loss function used to train the network is a weighted sum of the losses from all the iterations, with higher weights assigned to later iterations. Despite promising results, this approach exhibits lower performance compared to stacking modules. This may be attributed to the lack of specialized refinement modules, since the same network needs to leverage both relatively local information for initial segmentation and more structural and global information for further refinement. This limits the ability of the model to address the specific challenges of progressive segmentation refinement.
@@ 4 Methods
The Base subnetwork takes as input a retinography image and produces base segmentation maps of arteries (A), veins (V), and blood vessels (BV) (the union of arteries and veins). These segmentation maps (without the input image) are then fed to the RR subnetwork. This subnetwork recursively refines the segmentation maps of arteries and veins a certain number of iterations $K$, iteratively correcting manifest vessel classification errors made by the Base subnetwork. BV segmentation is not refined based on previous work indicating high accuracy with a single U-Net Morano et al., (2021); Karlsson and Hardarson, (2022). In each iteration, the input of the RR subnetwork is exclusively the output of the preceding iteration. This forces the network to focus on correcting errors based on the existing blood vessel structure (whose segmentation remains fixed through the different iterations), rather than relying on the characteristics of the input image. The final output of the network consists of the refined A/V segmentation maps at the last iteration ($k=K$) and the initial BV segmentation map produced by the Base subnetwork.
@@
where $w_{k}$ is the weight of the loss at iteration $k$, and $\mathcal{L}_{\text{S}}$ is the segmentation loss function, defined as the sum of individual binary segmentation errors for each structure (arteries, veins, and BV). Following previous works Morano et al., (2021); Chen et al., (2022), we use the Binary Cross-Entropy (BCE) loss between the prediction and the GT for each structure. Thus, the segmentation loss $\mathcal{L}_{\text{S}}$ for $N$ structures is defined as:
@@
where $\mathbf{\hat{y}}_{i}$ and $\mathbf{y}_{i}$ are the $i$-th channel of the output of the network and the GT, respectively, representing individual structures. The weighting scheme is similar to that of Mookiah et al., (2021), with the difference that we give a higher weight to the error of the first iteration ($k=0$), which solely relies on the Base subnetwork. This prioritizes producing fairly accurate base segmentations before applying subsequent refinements. Specifically, the weight $w_{k}$ at iteration $k$ is defined as
@@
The proposed network architecture consists of two nearly-identical encoder-decoder subnetworks connected in series ($\text{FCNN}_{1}$ and $\text{FCNN}_{2}$, in Fig. 3). While architecturally similar, the subnetworks differ in their function and the number of output channels. The first subnetwork, used to obtain the base segmentation maps, has 3 output channels (arteries, veins, and BV), and the second, used for iterative refinement, has 2 (arteries and veins). Similarly to the state of the art Xu et al., (2018); Girard et al., (2019); Hemelings et al., (2019); Galdran et al., (2019); Ma et al., (2019); Morano et al., (2020, 2021), we adopt the original U-Net architecture Ronneberger et al., (2015) for both subnetworks. At the end of each subnetwork, a sigmoid function is used to produce the segmentation maps of all structures. A complete diagram of the U-Net architecture used in this work is shown in Fig. 4.

### v3

@@ 1 Introduction
The characteristics of retinal blood vessels (BV), including their caliber and configuration, serve as valuable biomarkers for diagnosing and monitoring several diseases and medical conditions, such as glaucoma, age-related macular degeneration (AMD), diabetic retinopathy (DR), and hypertension Sun et al., [2009], Abràmoff et al., [2010], Kanski and Bowling, [2011]. These alterations can be readily identified by trained ophthalmologists through analysis of color fundus images acquired via retinography, a non-invasive and cost-effective imaging technique that involves capturing photographs of the retina through the dilated pupil. By virtue of its affordability and lack of invasiveness, retinography has become widely adopted in clinical practice, research investigations, and population-wide screening programs. An example of a retinography image is shown in Fig. 1 (left).
@@
A comprehensive analysis of the retinal vasculature requires the segmentation of blood vessels and their subsequent classification into arteries and veins (A/V). This process yields separate A/V segmentation maps (as illustrated in Fig. 1, right), enabling the quantification of various diagnostically relevant vessel characteristics, such as width, diameter, and tortuosity. Furthermore, accurate measurement of these characteristics facilitates the calculation of more complex biomarkers, including the arteriolar-to-venular diameter ratio (AVR) Ikram et al., [2004], Sun et al., [2009], Hatanaka et al., [2005].
However, manual execution of these tasks is inherently laborious, leading to increased costs, and is prone to human error, which negatively impacts both reproducibility and quality of care. To address these limitations, several automated approaches have been proposed for the simultaneous segmentation and classification of arteries and veins Mookiah et al., [2021].
Current state-of-the-art methods predominantly leverage fully convolutional neural networks (FCNNs) Long et al., [2015] for this purpose Galdran et al., [2019], Hemelings et al., [2019], Morano et al., [2021], Chen et al., [2022], Galdran et al., [2022], Karlsson and Hardarson, [2022], Hu et al., [2024]. Most prevalent approaches classify each pixel into one of four classes: background, artery, vein, and crossing (representing regions where arteries and veins overlap). Additionally, some methods incorporate an “uncertain” class to account for pixels presenting ambiguous characteristics Galdran et al., [2019], Hemelings et al., [2019], Morano et al., [2021], Chen et al., [2022], Galdran et al., [2022], Karlsson and Hardarson, [2022], Hu et al., [2024]. Conversely, some recent approaches Chen et al., [2022], Morano et al., [2021] frame the A/V segmentation and classification task as a multi-label segmentation problem. This framework entails training the network to simultaneously segment arteries, veins, and BV (i.e., both arteries and veins) as separate classes, allowing a single pixel to be assigned to one or more classes.
@@
Figure 2: Examples of manifest classification errors produced by a state-of-the-art FCNN-based method Morano et al., [2021]. (1-3) While most of the vessel is classified as artery, the model misclassifies the last part as vein. (4-6) The model often confuses the classification of vessels in crossing areas. These errors are easily detected by a human observer because they are inconsistent with the overall structure of the vascular tree, hence the term “manifest”.
These errors arise from the propensity of FCNN-based models to classify vessels based on local characteristics of the input image, neglecting the global structural context of the vascular tree. To mitigate these errors, several approaches have been proposed Girard et al., [2019], Kang et al., [2020], Chen et al., [2022], Hu et al., [2024]. Some methods Girard et al., [2019], Kang et al., [2020] employ ad hoc post-processing techniques based on graph propagation. Alternatively, other methods Chen et al., [2022], Hu et al., [2024] have proposed combining standard pixel-wise segmentation losses with adversarial losses Goodfellow et al., [2014] and custom-designed losses focused on specific characteristics of the predicted maps, such as topological consistency. This combined approach aims to guide the model towards generating more topologically consistent segmentations. Despite their contributions, these approaches exhibit limitations in their applicability beyond the specific task of blood vessel segmentation and classification (i.e., limited generalizability), and their overall effectiveness in mitigating manifest classification errors remains limited.
@@ 2.1 Vessel segmentation and classification
The first methods for vessel segmentation on color fundus images were based on ad hoc image processing techniques Jiang and Mojon, [2003], Nain et al., [2004], Tolias and Panas, [1998], Staal et al., [2004] or traditional learning models such as artificial neural networks Sinthanayothin et al., [1999], Marín et al., [2011]. Today, FCNNs based on the U-Net architecture Ronneberger et al., [2015] have become the state of the art Jiang et al., [2018], Oliveira et al., [2018], Jin et al., [2019], Wang et al., [2021], Liu et al., 2023c , Liu et al., 2023a , Liu et al., 2023b .
Until recently, A/V segmentation and classification was treated as a two-step process, where A/V classification was performed only on pixels previously identified as BV by a segmentation algorithm Relan et al., [2014], Dashtbozorg et al., [2014], Zamperini et al., [2012], Estrada et al., [2015], Welikala et al., [2017]. In addition, many works restricted the classification to small regions, usually around the optic disc Zamperini et al., [2012], Relan et al., [2014]. The first work that deals with the whole vascular tree Dashtbozorg et al., [2014] proposed a graph-based method that takes as input a previously segmented vascular graph and the input image and obtains two separate graphs for arteries and veins. Later, Welikala et al., [2017] were the first to propose the use of a CNN for the classification stage. Although these methods achieved reasonable performance, they were limited by the quality of the initial vessel segmentation.
To avoid this problem, several works have addressed the simultaneous segmentation and classification of retinal vessels either as a semantic segmentation task of three to four classes (background, artery, vein, uncertain) Xu et al., [2018], Girard et al., [2019], Hemelings et al., [2019], Galdran et al., [2019], Ma et al., [2019], Kang et al., [2020], Morano et al., [2021], Chen et al., [2022], Karlsson and Hardarson, [2022], Galdran et al., [2022] or as a multi-label segmentation task with multiple targets Morano et al., [2021], Chen et al., [2022] (arteries, veins, blood vessels). The latter approach has the advantage of providing continuous and thus more topologically consistent segmentation maps of arteries and veins, since vessel crossings are considered as both classes at the same time.
@@ 2.2 Reducing manifest classification errors
Ad hoc post-processing methods Girard et al., [2019], Kang et al., [2020] typically involve graph-based operations on the vessel graph extracted from previously segmented maps. These methods usually require additional input information (such as the location of the optic disc), and their effectiveness is severely limited by the quality of the initial segmentation maps, which hinders their generalization capabilities.
Conversely, learning-based methods Hu et al., [2024], Chen et al., [2022], Karlsson and Hardarson, [2022] aim to address manifest classification errors by incorporating additional losses and architectural modifications. Hu et al., [2024] introduced a multi-class point consistency module to generate artery and vein skeletons, which are then used to compute different consistency losses aimed at improving topological consistency and mitigating classification errors. Similarly, Chen et al., [2022] proposed a GAN-based method with a topological loss. In particular, they proposed to use a discriminator designed to rank, from lowest to highest, the topological connectivity of the ground truth, the predicted mask, and a randomly transformed mask. The ranking error is used as a loss to encourage the model to produce more topologically consistent segmentation maps. In addition, they proposed a new module that extracts the high-level topological features of the images to force the model to predict vascular segmentation maps with a topology similar to that of the manual annotations. Although these methods are interesting and fairly effective, they rely on task-specific mechanisms, and their performance remains limited.
@@ 2.3 Iterative refinement
In recent years, several iterative refinement approaches have been proposed to improve segmentation performance in both natural and medical image analysis Pinheiro and Collobert, [2014], Sironi et al., [2016], Newell et al., [2016], Shen et al., [2017], Januszewski et al., [2016], Mosinska et al., [2018], Karlsson and Hardarson, [2022], Galdran et al., [2022]. These methods use an iterative prediction process via a classifier that receives as input the result from the previous iteration(s) and, optionally, the input image, addressing the errors made in earlier iterations.
@@
A common approach consists of stacking multiple deep modules and training them in an end-to-end fashion Newell et al., [2016], Shen et al., [2017], Karlsson and Hardarson, [2022], Galdran et al., [2022]. This allows for modules specialized in solving the errors of the previous modules. For example, Newell et al., [2016] proposed a novel architecture composed of eight consecutive modules for pose estimation in natural images. During training, all modules receive supervision through comparison of their outputs with the ground truth. However, such methods often require a substantial number of parameters, leading to high memory and computational costs during both training and inference. To address this limitation, recent work has proposed the use of very lightweight encoder-decoder networks Karlsson and Hardarson, [2022], Galdran et al., [2022]. Using a custom architecture consisting of four U-Net-like networks, Karlsson and Hardarson, [2022] achieved state-of-the-art performance on various A/V segmentation and classification benchmarks. However, their approach requires extensive hyperparameter tuning to perform well, and the optimal hyperparameters were found to be dataset-specific. Their final model was achieved through an exhaustive search exploring various network configurations and loss functions. In particular, the authors experimented with different numbers of networks, layers, levels, and kernels in each network/layer, as well as different weights for the loss terms and the regularization parameters. Additionally, the generalization capabilities of the method have not been thoroughly evaluated. Galdran et al., [2022] proposed a similar approach for the same task using only two stacked U-Net-like networks, although with limited performance.
@@
An alternative approach consists of refining the predictions using a single recursive network Pinheiro and Collobert, [2014], Mosinska et al., [2018]. While the memory requirements during training of this approach remain constant, as the gradients for each iteration must be stored, the number of parameters is much lower, which makes it more efficient at test time. In this line of work, Pinheiro and Collobert, [2014] proposed to perform semantic segmentation of natural images by using a CNN network that subsequently refines the predicted maps at different scales. This approach is focused on increasing the spatial context of the network, so that it models non-local dependencies (of higher level) in the scenes. In this way, the authors manage to make the network give rise to more structurally coherent predictions. The problem with this method is that the CNN is applied pixel-wise, making it very inefficient in terms of computational cost. This problem is addressed by Mosinska et al., [2018] by using a FCNN at a constant scale. Their method involves training a FCNN that recursively refines an initial segmentation over multiple iterations. In the first iteration, the network receives the input image and an empty segmentation map (all zeros). In subsequent iterations $k\in\{1,...,K\}$, it receives the same image together with the segmentation map obtained in iteration $k-1$. The training loss function used to train the network is a weighted sum of the losses from all the iterations, with higher weights assigned to later iterations. Despite promising results, this approach exhibits lower performance compared to stacking modules. This may be attributed to the lack of specialized refinement modules, since the same network needs to leverage both relatively local information for initial segmentation and more structural and global information for further refinement. This limits the ability of the model to address the specific challenges of progressive segmentation refinement.
@@ 4 Methods
The Base subnetwork takes as input a retinography image and produces base segmentation maps of arteries (A), veins (V), and blood vessels (BV) (the union of arteries and veins). These segmentation maps (without the input image) are then fed to the RR subnetwork. This subnetwork recursively refines the segmentation maps of arteries and veins a certain number of iterations $K$, iteratively correcting manifest vessel classification errors made by the Base subnetwork. BV segmentation is not refined based on previous work indicating high accuracy with a single U-Net Morano et al., [2021], Karlsson and Hardarson, [2022]. In each iteration, the input of the RR subnetwork is exclusively the output of the preceding iteration. This forces the network to focus on correcting errors based on the existing blood vessel structure (whose segmentation remains fixed through the different iterations), rather than relying on the characteristics of the input image. The final output of the network consists of the refined A/V segmentation maps at the last iteration ($k=K$) and the initial BV segmentation map produced by the Base subnetwork.
@@
where $w_{k}$ is the weight of the loss at iteration $k$, and $\mathcal{L}_{\text{S}}$ is the segmentation loss function, defined as the sum of individual binary segmentation errors for each structure (arteries, veins, and BV). Following previous works Morano et al., [2021], Chen et al., [2022], we use the Binary Cross-Entropy (BCE) loss between the prediction and the GT for each structure. Thus, the segmentation loss $\mathcal{L}_{\text{S}}$ for $N$ structures is defined as:
@@
where $\mathbf{\hat{y}}_{i}$ and $\mathbf{y}_{i}$ are the $i$-th channel of the output of the network and the GT, respectively, representing individual structures. The weighting scheme is similar to that of Mookiah et al., [2021], with the difference that we give a higher weight to the error of the first iteration ($k=0$), which solely relies on the Base subnetwork. This prioritizes producing fairly accurate base segmentations before applying subsequent refinements. Specifically, the weight $w_{k}$ at iteration $k$ is defined as
@@
The proposed network architecture consists of two nearly-identical encoder-decoder subnetworks connected in series ($\text{FCNN}_{1}$ and $\text{FCNN}_{2}$, in Fig. 3). While architecturally similar, the subnetworks differ in their function and the number of output channels. The first subnetwork, used to obtain the base segmentation maps, has 3 output channels (arteries, veins, and BV), and the second, used for iterative refinement, has 2 (arteries and veins). Similarly to the state of the art Xu et al., [2018], Girard et al., [2019], Hemelings et al., [2019], Galdran et al., [2019], Ma et al., [2019], Morano et al., [2020, 2021], we adopt the original U-Net architecture Ronneberger et al., [2015] for both subnetworks. At the end of each subnetwork, a sigmoid function is used to produce the segmentation maps of all structures. A complete diagram of the U-Net architecture used in this work is shown in Fig. 4.